Session: Mar 2008 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Tao An at Thu Feb 28 14:03:34 2008) Success :

Medicina (A.Orlati at Thu Feb 28 14:18:48 2008) Success :

Torun (Kaz Borkowski at Thu Feb 28 14:57:38 2008) Success :
No known problems. Earlier Bob informed on a Tr problem of phase jumps in C band observations. We took certain measures, but we still do not know the source and whether it affects this L band.

Urumqi (Liu Xiang at Thu Feb 28 15:18:31 2008) Success :
the first 20min failed for a little problem, otherwise ok, it is cloudy and snow at last hour.

Onsala (Michael Lindqvist at Thu Feb 28 20:47:01 2008) Success :
No known problems.

Socorro (Tao An at Fri Feb 29 04:34:51 2008) Minor failures : RECO
1.the warning "recording throttled can't keep up " for everyscan. 2.the error "scan_check" failed for some scans.

Hartebeesthoek (Jonathan Quick at Fri Feb 29 07:28:52 2008) Success :
No known problems

Westerbork (Tony Foley at Fri Feb 29 13:38:14 2008) Minor failures :
Wb used in phased array mode. started slightly late

Effelsberg (Uwe Bach at Mon Mar 3 12:02:19 2008) Minor failures : HIGH
Started observations with a warm receiver. Cooling went down from 90K to 50K during observations. Nominal Tsys should be about 35K.

Jodrell_Bank (Alastair Gunn at Mon Mar 3 15:43:35 2008) Success :
The Lovell Telescope was used (Jb1). No known problems.

Cambridge (Alastair Gunn at Mon Mar 3 15:44:00 2008) Success :
No known problems.

Noto (P.Cassaro at Tue Mar 4 16:52:07 2008) Success :